A survey done of 25 voters revealed that 17 voters where in favor of a recent change and 8 opposed it. If 6 random voters where chosen from the survey what is the probability that exactly 2 voters where in favor of the change, and exactly 4 opposed it?
Answer by stanbon(75887)

A survey done of 25 voters revealed that 17 voters where in favor of a recent change and 8 opposed it. If 6 random voters where chosen from the survey what is the probability that exactly 2 voters were in favor of the change, and exactly 4 opposed it?
-----------------
It's binomial with n = 6 ; p = 17/25 ; x = 2
---
P(x=2) = 6C2(17/25)^2(8/25)^4 = binompdf(6,17/25,2) = 0.0727...
Cheers,
Stan H.